OpenStudy (jacobbenvenutty):

What is the equation in standard form of the line which passes through (−2, −3) and has a slope of one half? Please don't just give me the answer, I would like to know how to do it for future problems.

OpenStudy (anonymous):

The standard form of a line looks like this: Ax+By=C Since you are given the slope of the line as 1/2, I would use the slope point form of a line to find the constants in the standard form of a line. The slope form of a line looks like this: y=mx+b. Where m is the slope of the line and small case b is where the line crosses the y axis. So substitute the values for x, y and slope in to the point slope equation and solve for b. -3 = 1/2(-2) +b -3 = -1 +b -2 = b Now just plug in the b value and slope into the point slope equation. Don't plug in the given values for the x and y. It should look like this: y = 1/2x -2 Now rearrange the equation. 1/2x -y = 2 Clear the fraction 2(1/2x) - 2(y) = 2(2) Standard form is: x-2y=4

OpenStudy (jacobbenvenutty):

Thank you very much